### v2

# SDSC:A Structure-Aware Metric for Semantic Signal Representation Learning

Abstract
We propose the Signal Dice Similarity Coefficient (SDSC), a novel structure-aware metric function for time series self-supervised representation learning. Most Self-Supervised Learning (SSL) methods for signals commonly adopt distance-based objectives such as mean squared error (MSE), which are sensitive to amplitude, invariant to waveform polarity, and unbounded in scale. These properties hinder semantic alignment and reduce interpretability. To the best of our knowledge, SDSC is the first similarity metric explicitly designed to capture structural alignment in temporal signals. SDSC addresses this by quantifying structural agreement between temporal signals based on the intersection of signed amplitudes, derived from the Dice Similarity Coefficient (DSC).Although SDSC is defined as a structure-aware metric, it can be used as a loss by subtracting from 1 and applying a differentiable approximation of the Heaviside function for gradient-based optimization. A hybrid loss formulation is also proposed to combine SDSC with MSE, improving stability and preserving amplitude where necessary. Experiments on forecasting and classification benchmarks demonstrate that SDSC-based pre-training achieves comparable or improved performance over MSE, particularly in in-domain and low-resource scenarios. The results suggest that structural fidelity in signal representations enhances the semantic representation quality, supporting the reconsideration of structure-aware metrics as viable alternatives to conventional distance-based methods.
Introduction
SSL enables representation learning from unlabeled data by formulating proxy objectives such as masked prediction or reconstruction. In Computer Vision (CV) and Natural Language Processing (NLP), SSL methods have demonstrated effective feature extraction capabilities. (Gui et al. 2024)
In time series modeling, SSL is also applied to extract semantic representations for downstream tasks, including forecasting and classification. (Ma et al. 2024) In time-series signals such as EEG and EMG, task-relevant semantics are often encoded in structural features, including waveform shapes, phase alignment, and local frequency patterns. Signal-specific representations require attention to these temporal structures. However, most of the reconstruction-based SSL methods commonly adopt distance-based metrics such as MSE, which focus on amplitude differences and element-wise deviations. Distance-based metrics exhibit sensitivity to scale and are invariant to waveform polarity, resulting in reconstructions that minimize error without preserving structural consistency. In addition, distance-based metrics often assign low values to reconstructions that differ semantically from the target. Phase-inverted signals, amplitude-scaled signals, and zero-valued baselines produce similar MSE values despite large semantic deviations. Moreover, metric values are unbounded and non-normalized, which complicates interpretation and model selection. Therefore, the lack of interpretability and amplitude bias in MSE limit its reliability.
To address these issues, a novel structure-aware metric, SDSC is introduced. The SDSC is inspired by the DSC (Dice 1945; Sørensen 1948) widely used in semantic segmentation tasks and quantifies the structural agreement between two temporal signals. The metric is bounded in [0, 1], reduces sensitivity to amplitude variation, and reflects alignment in waveform structure. In addition, the metric is implemented as a loss by subtracting values from 1 and a differentiable approximation of the Heaviside function is applied to enable gradient-based optimization. Lastly, a hybrid loss that combines SDSC and MSE is introduced to enhance stability during training. Experiments on forecasting and classification benchmarks show that SDSC-based pre-training achieves comparable or improved performance relative to MSE, particularly when encoders are fixed. Performance gains are observed in in-domain and low-resource settings, where structural fidelity in the representations contributes to downstream accuracy.
Self-supervised learning for time series signals typically consists of two components: reconstruction-based and contrastive learning. While contrastive objectives explicitly enforce instance discrimination, reconstruction-based objectives aim to recover the signal content. In this context, the choice of reconstruction loss directly influences the quality of the learned representation. Despite their widespread use, distance-based losses do not reflect the semantic structure and amplify the sensitivity to signal amplitude. The empirical similarity in downstream performance between MSE and SDSC indicates that MSE-based models achieve competitive results not due to accurate semantic preservation but due to incidental alignment with signal structure. These observations indicate that distance-based metrics might not fully capture the semantic structure of signals, motivating consideration of a shift towards structure-aware representation learning paradigms.
Related Work
This paper is related to DSC (Dice 1945; Sørensen 1948), and the broader field of time series modeling (TSM), specifically time series pre-training modelings (TS-PTM).
Evaluation Metrics
Traditional TS-PTM research widely used reconstruction metrics such as MSE or MAE. However, these metrics do not consider temporal misalignments between signals, which present significant limitations for analyzing various time series data. To overcome this, Dynamic Time Warping (DTW) was proposed in (Sankoff and Kruskal 1983), effectively addressing misalignment issues. Despite its effectiveness, DTW suffers from computational inefficiency. To mitigate this, (Salvador and Chan 2007) introduced Fast DTW, significantly reducing computation time but remaining unsuitable for training due to its non-differentiability. Soft DTW(Cuturi and Blondel 2017) addressed this limitation by providing a differentiable approximation suitable for training. However, these metrics are fundamentally distance-based.
The Dice Score Coefficient (DSC) (Dice 1945; Sørensen 1948) is commonly used in CV, especially semantic segmentation. This paper investigates applying DSC to time series rather than computer vision datasets. Specifically, our proposed SDSC adapts DSC for time series, enabling overlap-based evaluations sensitive to signal structure rather than amplitude.
Time Series Modeling
According to (Gui et al. 2024), SSL has advanced in various domains, typically involving a two-stage process of unsupervised pre-training followed by task-specific fine-tuning. SSL has notably improved performance across diverse time series tasks, becoming a critical step in contemporary research.
Several notable TS-PTM studies have sought to enhance performance through various strategies (Ma et al. 2024). TS2Vec (Yue et al. 2022) proposed an architecture for contextual representations. CoST (Woo et al. 2022) adopted contrastive learning to improve TS-PTM. TmesNet (Wu et al. 2022) addressed general time series analysis and (Eldele et al. 2023) investigated semi-supervised time series learning using contrastive methods. TI-MAE (Li et al. 2023) leveraged masked autoencoders for improved representation learning, and SimMTM (Dong et al. 2023) further simplified the masked TS-PTM frameworks effectively. In addition, iTransformer (Liu et al. 2023) introduced inverted transformers for efficient forecasting, and Unified Transformer (Woo et al. 2024) generalized a single training approach to multiple downstream tasks. TIMER (Liu et al. 2024) developed generative pre-trained transformers for large-scale time series modeling. Furthermore, TimeSiam (Dong et al. 2024) proposed a Siamese network-based pre-training framework, and TimeDiT (Cao et al. 2024) suggested diffusion-based TS-PTM. Cross-domain research on large-scale data was conducted by (Prabhakar Kamarthi and Prakash 2024), while (He et al. 2025) investigated universal knowledge representation for general TS-PTM.
Although these works contributed significantly to the progress of TSM, most focused on architectural innovations or adopted contrastive strategies, which can be limiting to learning structural representations. To the best of our knowledge, no prior work has proposed a structure-aware metric for signal representation learning. To address this, we introduce a structure-aware metric tailored to time-series data. SimMTM (Dong et al. 2023) is selected as our experimental baseline for its simplicity, strong empirical performance, and consistent application of traditional metrics, providing a reliable foundation to assess the impact of our proposed SDSC.
Signal Dice Similarity Coefficient
In this section, we introduce the SDSC, a new metric designed to explicitly quantifies structural similarities between two signals. In signal representation learning, reconstructing signals accurately is important to capture their meaning.
Figure 1: Examples demonstrating the limitations of distance-based metrics in capturing structural similarity. SDSC offers a more faithful assessment in (top-left) phase-shifted signals, (top-right) scale-induced distortions, (bottom-left) structurally dissimilar but MSE-equivalent signals, and (bottom-right) noisy outputs with underestimated errors.
Signals MSE↓ MAE↓ DTW↓ SDSC↑
Inverted 0.0200 0.1272 0.0425 0.0000
0.5x Scaled 0.1249 0.3180 0.1353 0.6667
2x Scaled 0.4995 0.6360 0.2706 0.6667
Zero 0.4995 0.6360 0.6360 0.0000
Noise Sample 0.5062 0.6361 0.2236 0.1137
Positive Shifted 1.0000 1.0000 0.6228 0.3887
Negative Shifted 1.0000 1.0000 0.6228 0.3887
Table 1: Quantitative evaluation of signal variations using distance-based metrics and the proposed SDSC
Distance-based Metrics
Distance-based metrics such as MSE, mean absolute error (MAE) and dynamic time warping (DTW) are widely used to measure the difference between predicted and ground-truth signals. These metrics evaluate element-wise deviations and are effective in reducing numerical reconstruction errors. However, distance-based metrics focus primarily on signal amplitude and do not consider the polarity or structural shape of the waveform.
Figure 1 and Table 1 illustrate examples that expose key limitations of distance-based metrics. Figure 1 (top left) illustrates a complete phase inversion under low-amplitude conditions, visually preserving the waveform shape while reversing its polarity. As shown in Table 1 (Inverted), the inverted signal receives low error scores on all distance-based metrics (for example, MSE = 0.0200), making it appear as a high quality reconstruction despite its semantic reversal. Figure 1 (top right) compares 0.5× and 2× scaled signals, both of which introduce comparable structural distortions but produce markedly different metric values due to amplitude differences. As indicated in Table 1, the dependence on the amplitude obscures the true degree of structural deviation, resulting in an inaccurate evaluation of the signal quality. In Figure 1 (bottom left), a constant zero signal is evaluated alongside a 2× scaled waveform. As shown in Table 1, both produce identical MSE scores (0.4995), despite their stark structural differences. The similarity in MSE scores despite structural differences reveals the inability of MSE to distinguish between waveforms when average magnitudes are equivalent. Lastly, Figure 1 (bottom right) contrasts noisy outputs with semantically valid yet shifted waveforms. The noise-dominated signal produces an MSE (0.5062) comparable to semantically valid signals due to its fluctuation averaging around the baseline, making it appear deceptively accurate under distance-based metrics. Although numerically favorable, the output is structurally misaligned and functionally misleading. Such insensitivity to signal semantics is particularly problematic for physiological data like EEG or ECG, where subtle structural components often carry diagnostic significance. Therefore, exclusive reliance on amplitude-centric metrics may lead to semantically incorrect reconstructions.
Definition of SDSC
The SDSC extends the DSC, commonly used for set overlap in semantic segmentation, to continuous, signed time-series data. Instead of relying on the membership of the sets, SDSC computes structural alignment from signed amplitude intersections at each time step. This formulation captures polarity, phase, and alignment—factors often ignored by distance-based metrics like MSE. Like DSC, SDSC returns a score in the range $[0,1]$. The original DSC measures set the similarity as follows:
$DSC=\frac{2|X\cap Y|}{|X|+|Y|}$ (1)
Here, $|X|$ and $|Y|$ denote the cardinalities of the respective sets, and the metric reflects the size of their intersection relative to the total area. The SDSC extends this concept to the signal domain by interpreting the area under the curve as a proxy for the waveform structure. Given two signal functions $E(t)$ and $R(t)$ representing ground truth and reconstruction, the SDSC is defined as :
$S(t)=E(t)\cdot R(t)$ (2)
$M(t)=\frac{(|E(t)|+|R(t)|)-(|E(t)|-|R(t)|)}{2}$ (3)
[…]
$H(\cdot)$ denotes the Heaviside step function, and $t\in T$ is given time. The objective in signal representation learning is to maximize Equation (4) toward 1. However, directly computing SDSC via integration is infeasible in practice, as real-world signals, such as EEG, lack known analytical expressions. To address this, a discrete approximation is adopted.
Figure 2: Example of intersection between two signals and discrete approximation.
Figure 2 illustrates the approximation procedure. Although signals are continuous in nature, real-world signals are typically sampled at uniform intervals. Consequently, each sampled value is treated as a rectangle of unit width, allowing the continuous integral to be approximated by summation.
$SDSC(E(t),R(t))\approx\frac{2\cdot\sum H(S(s))\cdot M(s)}{\sum(E(s)+R(s))}$ (5)
where $s\in S$ are discrete sampling points, with $S\subset T$. The proposed approximation enables a tractable computation of SDSC on real signals and ensures a consistent evaluation. As demonstrated in Table 1, the SDSC shows increased robustness to polarity shifts and amplitude scaling. The output range [0,1] facilitates standardized interpretation in all signal domains. Unlike MSE or DTW, the SDSC is less affected by signal magnitude, thereby reducing distortions due to scale and polarity. The normalized form also simplifies cross-domain comparisons, enabling a more structure-aware assessment of signal reconstruction quality.
Hybrid Loss Integration
Since the SDSC score is bounded in $[0,1]$, we can define the loss as $1-SDSC(\cdot)$.
$\mathcal{L}_{sdsc}=1-SDSC(E(S),R(S))$ (6)
However, the use of the Heaviside step function in Equation (4) introduces discontinuities, which can negatively affect the stability of training. Continuity is preserved when at least one of the signals maintains the same sign at the corresponding sampled points. However, the likelihood of sign mismatches increases when the sampling resolution is low. To enable stable gradient-based optimization, a smooth approximation of the Heaviside function is introduced. The following sigmoid-based formulation is used, with a sharpness parameter $\alpha$.
$\hat{H}(x)=\frac{1}{1+e^{-\alpha x}}\quad$ (7)
If the sharpness parameter $\alpha$ is large, the sigmoid-based approximation $\hat{H}(x)$ more closely resembles the original Heaviside function. However, excessively large values of $\alpha$ can lead to sharp transitions that result in unstable gradients, potentially damaging the training process.
SDSC captures structure but ignores amplitude, whereas MSE captures amplitude but misses structure. To balance the strengths of both approaches, this work proposes a hybrid loss function that combines the structural awareness of SDSC with the amplitude sensitivity of MSE. The final objective function is formulated as follows:
$\mathcal{L}_{hybrid}=\lambda_{sdsc}\cdot\mathcal{L}_{sdsc}+\lambda_{mse}\cdot\mathcal{L}_{MSE}$ (8)
Here, $\lambda_{sdsc}$ and $\lambda_{mse}$ are parameters that control the trade-off between structural accuracy and amplitude-based accuracy. To determine these weights, we adopt the uncertainty-based tuning strategy proposed in  (Kendall, Gal, and Cipolla 2018), where the weighting coefficients are adapted based on the homoscedastic uncertainty associated with each loss term. This hybrid formulation promotes reconstructions that are structurally aligned and numerically precise.
Experiments
All experiments are conducted with fixed random seeds across all runs to ensure reproducibility. Training and evaluation are performed on NVIDIA 3090 GPUs (2×). For fair comparison, we adopt SimMTM (Dong et al. 2023) as the sole baseline model throughout all experiments. Baseline models are reproduced using their official implementations, and all hyperparameters follow the original settings reported in their respective papers. Due to resource limitations, the input sequence length is fixed at 96 for forecasting experiments.
For time series forecasting tasks, we adopt MSE and MAE as evaluation metrics. For classification experiments, accuracy, precision, recall, and F1 score are computed, along with their macro-average values.
Gradient Sensitivity Analysis
In this section, we measure the gradient norm with respect to different types of signal perturbations to assess how each loss function responds to structural variations. The analysis is based on the same toy cases as in Table 1, including the jitter case.
  Example MSE MAE SDSC
Inverted 0.0894 0.0316 0.0
0.5x Scaled 0.0223 0.0316 0.0442
2x Scaled 0.0447 0.0316 0.0110
Zero 0.0447 0.0316 0.0
Noise Sample 0.0194 0.0316 0.0237
Shifted 0.0632 0.0316 0.0075
Jittered 0.0032 0.0316 0.0248
Table 2: Gradient Sensitivity Table
  Example $\alpha=1$ $\alpha=10$ $\alpha=100$
Inverted 0.0091 0.0082 0.0047
0.5x Scaled 0.0289 0.0437 0.0436
2x Scaled 0.0062 0.0102 0.0102
Zero 0.0 0.0 0.0
Noise Sample 0.0152 0.0228 0.0237
Shifted 0.0074 0.0087 0.0076
Jittered 0.0165 0.0228 0.0242
Table 3: Gradient Sensitivity with alpha Table
Tables 2 and 3 present the gradient norms for MSE, MAE, and SDSC, and the effect of varying the $\alpha$ parameter in SDSC, respectively.
MSE exhibits significant gradient changes under amplitude perturbations, but it fails to respond meaningfully to minor structural variations, such as jitter. In contrast, SDSC yields low gradients for structure-preserving signals (e.g., shifted), while responding with larger gradients when local waveform patterns are distorted (e.g., jittered). Although gradient vanishing may occur in some structure-breaking cases, SDSC remains robust to amplitude shifts.
As $\alpha$ increases, the behavior of the approximated SDSC becomes more similar to that of the original formulation. The results suggest that $\alpha=10$ is sufficient for a close approximation in practice.
Losses MSE SDSC Hybrid
Metrics MSE↓ MAE↓ SoftDTW↓ SDSC↑ MSE↓ MAE↓ SoftDTW↓ SDSC↑ MSE↓ MAE↓ SoftDTW↓ SDSC↑
ETTh1 1.1980 0.6863 5780 0.5573 1.5430 0.7653 6703 0.5725 1.2070 0.6738 5425 0.5774
ETTh2 0.6229 0.4595 2616 0.7263 0.6553 0.4643 2688 0.7206 0.6210 0.4487 2603 0.7320
ETTm1 1.1740 0.6532 5652 0.5759 1.7880 0.7975 7738 0.5838 1.1780 0.6420 5570 0.5886
ETTm2 0.0269 0.1121 -986.2 0.9355 0.0280 0.1138 -978.2 0.9345 0.0242 0.1052 -996.4 0.9395
Weather 0.0863 0.1126 253.9 0.9022 0.1024 0.1045 483 0.9097 0.0806 0.1007 163.3 0.9132
Electricity 0.0970 0.2136 11937 0.8544 0.1039 0.2215 12794 0.8524 0.0842 0.1960 10366 0.8679
Traffic 0.1914 0.2304 15723 0.8382 0.2229 0.2418 18217 0.8327 0.1529 0.1914 12600 0.8700
Avg 0.4852 0.3525 5953.6714 0.7670 0.6348 0.3870 6806.4 0.7723 0.4783 0.3368 5104.4143 0.7841
Epilepsy 0.0807 0.2456 1839 0.6244 0.1300 0.2740 2962 0.6736 0.1099 0.2484 2503 0.6856
SleepEEG 100.5599 6.8082 2291156 0.5966 147.9205 7.4512 3370220 0.6483 100.5852 6.8087 2291733 0.6105
Avg 50.3203 3.5269 1146498 0.6105 74.0253 3.8626 1686591 0.6610 50.3471 3.5286 1147118 0.6481
Table 4: Pretraining Example
PreTraining
In the pre-training phase, we compare three objective functions for SSL: MSE loss, SDSC loss, and a hybrid loss that combines both. Table 4 summarizes the results in both the forecasting and classification datasets.
MSE-based models achieve lower reconstruction errors under distance-based metrics, as expected. In contrast, SDSC-based models achieve higher SDSC scores, which are indicative of structural alignment. While the metrics are weakly correlated, each captures distinct aspects of the signal, motivating a closer examination of their downstream impact.
Figure 3: Scatter plot between MSE and SDSC under MSE-based pretraining. Weak negative correlation suggests limited structural alignment.
Figure 4: SDSC distribution at MSE $=1.5\pm\epsilon$ for MSE-based (left) and SDSC-based (right) models. SDSC-based pretraining achieves higher structural scores.
  std IQR
MSE-based model 0.0280 0.0418
SDSC-based model 0.0249 0.0384
Table 5: Comparison of SDSC concentration under fixed MSE $(1.5\pm\epsilon)$.
Figure 3 shows the relationship between MSE and SDSC for models pre-trained on ETTh1 using MSE. SDSC increases as MSE decreases, but the Pearson correlation is $-0.324$, indicating a weak alignment. Figure 4 compares the SDSC distributions at fixed MSE ($1.5\pm\epsilon$) under MSE-based and SDSC-based pretraining. SDSC-based pretraining achieves higher SDSC values at the same MSE level. Table 5 reports the standard deviation and interquartile range (IQR) of SDSC values under fixed MSE. SDSC-based models exhibit lower variance and tighter concentration, suggesting a more consistent structural alignment.
These results indicate that MSE-based SSL captures structural features to some extent but lacks reliability. The weak correlation explains why SDSC values remain low when models are trained solely with MSE. SDSC-based training enhances structural fidelity but reduces amplitude precision, resulting in higher MSE. The hybrid loss addresses this trade-off and achieves stable performance across both metrics.
Scenarios Loss function Accuracy (%) Precision (%) Recall (%) F1 (%) Avg (%)
In Domain Epilepsy $\downarrow$ Epilepsy MSE 90.69 93.23 77.23 82.26 85.85
SDSC 91.86 93.82 80.27 84.98 87.73
Hybrid 92.44 93.58 82.13 86.38 88.63
SleepEEG $\downarrow$ SleepEEG MSE 60.20 52.90 49.75 46.92 52.44
SDSC 60.89 54.55 49.59 46.72 52.94
Hybrid 60.01 51.86 49.09 46.55 51.88
Cross Domain SleepEEG $\downarrow$ Epilepsy MSE 80.21 40.11 50.00 44.51 53.71
SDSC 80.21 40.11 50.00 44.51 53.71
Hybrid 80.21 40.11 50.00 44.51 53.71
SleepEEG $\downarrow$ FD-B MSE 52.19 35.80 38.21 33.87 40.02
SDSC 51.71 37.11 37.86 31.84 39.63
Hybrid 51.43 34.34 37.65 34.01 39.96
SleepEEG $\downarrow$ Gesture MSE 70.00 64.80 70.00 66.17 67.74
SDSC 68.33 66.34 68.33 63.89 66.72
Hybrid 70.00 66.09 70.00 66.17 68.07
SleepEEG $\downarrow$ EMG MSE 46.34 15.45 33.33 21.11 29.06
SDSC 46.34 15.45 33.33 21.11 29.06
Hybrid 46.34 15.45 33.33 21.11 29.06
Table 6: Freeze classification performance.
Forecasting
The forecasting task is performed in an in-domain setting using models pre-trained with MSE, SDSC, and hybrid objectives. For consistency, the evaluation is conducted at the best epoch for both pre-training and fine-tuning. We adopt MSE and MAE as evaluation metrics, following standard practice in forecasting tasks.
  Pre-Model MSE SDSC Hybrid
Metrics MSE↓ MAE↓ MSE↓ MAE↓ MSE↓ MAE↓
ETTh1 0.380 0.408 0.379 0.406 0.382 0.406
ETTh2 0.304 0.350 0.306 0.352 0.304 0.350
ETTm1 0.327 0.363 0.324 0.364 0.325 0.364
ETTm2 0.185 0.275 0.191 0.278 0.188 0.276
Weather 0.169 0.213 0.168 0.213 0.169 0.215
Electricity 0.200 0.291 0.200 0.293 0.198 0.288
Traffic 0.497 0.312 0.492 0.309 0.494 0.315
Avg 0.295 0.316 0.294 0.316 0.294 0.316
Table 7: Forecasting Finetuning Example
Table 7 summarizes the results of the forecasting task. As shown in Table 4, pre-trained MSE models learn SDSC to some extent, and pre-trained SDSC models, although the SDSC model has higher reconstruction errors, achieve similar downstream accuracy. These results indicate that beyond a certain threshold, additional MSE reduction achieves diminishing returns, suggesting limited benefits of focusing solely on amplitude rather than structural alignment. SDSC models achieve a similar accuracy with significantly higher MSE, implying that structural alignment alone suffices for downstream prediction.
Classification
The classification task is performed on both in-domain and cross-domain classification tasks using pre-trained encoders. To assess the effect of pretraining, we compare models with frozen encoders and those with end-to-end fine-tuning. All models are evaluated using accuracy, precision, recall, and the F1 score.
Scenarios Loss function Accuracy (%) Precision (%) Recall (%) F1 (%) Avg (%)
In Domain Epilepsy $\downarrow$ Epilepsy MSE 94.20 94.64 86.77 90.02 91.40
SDSC 93.97 94.46 86.16 89.56 91.04
Hybrid 93.91 94.59 85.89 89.42 90.95
SleepEEG $\downarrow$ SleepEEG MSE 65.11 57.10 55.70 52.16 57.52
SDSC 65.22 57.55 54.92 51.81 57.375
Hybrid 65.13 56.62 56.30 51.00 57.26
Cross Domain SleepEEG $\downarrow$ Epilepsy MSE 95.19 94.30 90.20 92.07 92.94
SDSC 92.03 94.33 80.48 85.27 88.03
Hybrid 95.19 94.30 90.20 92.07 92.94
SleepEEG $\downarrow$ FD-B MSE 63.88 69.36 72.98 69.98 69.05
SDSC 65.94 72.33 73.45 72.29 71.00
Hybrid 64.26 68.02 73.50 70.06 68.96
SleepEEG $\downarrow$ Gesture MSE 78.33 79.85 78.33 77.13 78.41
SDSC 80.00 80.21 80.00 79.32 79.88
Hybrid 78.33 81.20 78.33 76.15 78.50
SleepEEG $\downarrow$ EMG MSE 97.56 98.33 98.04 98.14 98.18
SDSC 95.12 96.83 91.37 93.62 94.24
Hybrid 95.12 96.83 91.37 93.62 94.24
Table 8: classification performance.
Table 6 reports the results with frozen encoders, while Table 8 summarizes the performance with fine-tuning. When encoders are frozen, pre-trained SDSC models consistently outperform others in in-domain settings. In other settings, performance differences depend on the characteristics of each dataset. For instance, the epilepsy dataset relies heavily on amplitude patterns, where pre-trained MSE models perform better. In contrast, the gesture dataset depends on the waveform structure, and SDSC models consistently achieve higher accuracy. These results highlight the need to select the reconstruction objective according to the properties of the signal. Pre-trained hybrid models provide a consistent choice across datasets, maintaining balanced performance. Lastly, SDSC models consistently achieve higher precision, suggesting that structural reconstruction contributes to more semantically meaningful representations in signal-based classification.
Conclusions
In this paper, we introduced the Signal Dice Similarity Coefficient (SDSC), a structure-aware metric for semantic representation learning in time series data. Unlike traditional distance-based metrics, SDSC is defined over a normalized range of $[0,1]$ and is robust to amplitude variations. Since SDSC is not directly differentiable, we proposed a smooth approximation to enable its use as a loss. We also identified limitations in amplitude-sensitive tasks, where SDSC may overlook critical signal properties. To address this, we proposed a hybrid loss that combines SDSC with MSE.
We integrated these loss functions into SimMT (Dong et al. 2023) and evaluated them on forecast and classification tasks. During pre-training, SDSC and MSE showed distinct behaviors, but downstream performance often converged, revealing weak correlation between the two metrics. This suggests that MSE-based models partially capture structural features without explicitly modeling them. However, we observed that excessive MSE minimization does not improve forecasting performance, indicating limited benefits from additional amplitude accuracy. In classification tasks, SDSC consistently improved performance in in-domain settings when encoders were frozen. This highlights SDSC’s ability to preserve semantic structure in pre-trained representations. Although end-to-end fine-tuning reduced this advantage in some datasets, SDSC remained competitive.
In conclusion, our results question the default reliance on MSE in signal pre-training and position SDSC as a promising metric for structure-aware learning in time series domains.Future work may investigate the integration of SDSC in other self-supervised frameworks and its effect on domain adaptation or cross-modal tasks. In addition, understanding the relationship between structural similarity and task-specific generalization remains an open question.
Acknowledgments